The so-called golden rectangle is said to be extremely pleasing visually and was used often by ancient Greek and Roman architects. The length of a golden rectangle is approximately 1.6 times the width. Find the dimensions of a golden rectangle if its area is 9000m^2.
Answer by checkley77(12844)

X*1.6X=9,000
1.6X^2=9,000
X^2=9,000/1.6
X^2=5625
X=SQRT5625
X=75 M FOR THE WIDTH.
1.6*75=120 M FOR THE LENGTH.
PROOF:
75*120
9,000=9,000